DeeDee04:

Triangle is translated 6 units up and 3 units left to create triangle A'B'C'. If vertex A is at (-1,2) and vertex B is at (1,5) then vertex A' is at a.(5,1) b. (-4,8) and vertex B' is at a.(-2,11) b.(-7,-1)

AngeI:

This question doesn't look complete?

Narad:

Take any point (x,y), then the translation is \[\left(\begin{matrix}x' \\ y'\end{matrix}\right) = \left(\begin{matrix}x-3 \\ y+6\end{matrix}\right)\]